Hi! I am having some issues with this equation, I usually can solve the ones that are not fractions, fractions are hard for me :( thank you
x/2-6=3+x/5

Answer by scott8148(6628)

not that wild about fractions, myself

multiplying by 10 (LCD) ___ 5x - 60 = 30 + 2x ___ 3x = 90